An urn contains nine balls of which three are red, four are blue and two are green. Three balls are drawn at
random without replacement from the urn. The probability that the three balls have different colours is -

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Total number of ways of drawing 3 balls out of 9 is 9c3=84
Number of ways of drawing so that balls are of different colours is (3c1)(4c1)(2c1)=24
Therefore probability is =24/84=2/7.
Thanks and Regards
Shaik Aasif
askIITians Faculty